# Spin structures and phase diagrams of the $\text{spin-}\frac{5}{2}$ triangular-lattice antiferromagnet Na2BaMn(PO4)2 under magnetic field

### v1

@@ V Data Availability
The neutron data collected at the ILL are available at Refs. 50, 51.
@@ VI Acknowledgments
We acknowledge help from N. Qureshi for Mag2POL and from O. Fabelo for the initial X-ray single crystal diffraction measurements performed at ILL. We thank B. Normand and A. M. Läuchli for discussions and comments. This work was supported by the Czech Science Foundation GAČR under the Junior Star Grant No. 21-24965M (MaMBA), by the Czech Ministry of Education Youth and Sports Grant No. LUABA24056 (BaCQuERel) and by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (Grant No. 205602). Crystals were grown and characterized in MGML (mgml.eu), which is supported within the program of Czech Research Infrastructures (project No. LM2023065).

### v2

@@ V Data Availability
The neutron data collected at the ILL are available at Refs. data_D23, data_D23b. The data treatment and figure reproduction workflow (notebook, scripts, and raw data inputs) are available on Figshare [data].
@@ VI Acknowledgments
We acknowledge help from N. Qureshi for Mag2POL and from O. Fabelo for the initial X-ray single crystal diffraction measurements performed at ILL. We thank B. Normand and A. M. Läuchli for discussions and comments. This work was supported by the Czech Science Foundation GAČR under the Junior Star Grant No. 21-24965M (MaMBA), by the Czech Ministry of Education Youth and Sports Grant No. LUABA24056 (BaCQuERel), and by the NCCR MARVEL, a National Centre of Competence in Research, funded by the Swiss National Science Foundation (Grant No. 205602). Crystals were grown and characterized in MGML (mgml.eu), which is supported within the program of Czech Research Infrastructures (project No. LM2023065).